6.A bank has borrowed $250 and has received $500 worth of deposits. The bank has given out loans for $650 and has reserves of 150, and the bank holds no securities or any other assets. The bank capital is:

Question options:

$50

$100

$500

7.Banks face interest-rate risk when interest rates change, because:

Question options:

lower interest rates always lead to lower net income for banks.

assets always have high interest rates and liabilities always have low interest rates, and banks are unwilling to change with market rates.

assets and liabilities may have different times to maturity, so they are affected differently by changes in market interest rates.
Answer #1

6.

Answer: 1st option; $50

Total assets = Loans + Reserve = 650 + 150 = $800

Total liabilities = Deposits + Borrowings = 500 + 250 = $750

Capital = Total assets – Total liabilities = 800 – 750 = $50

7.

Answer: 3rd option

The interest-rate risk arises because of different maturity dates of various assets and liabilities. If the rate is high, the interest payments to liabilities would be high; it decreases net income. If the rate is low, the immediate interest earnings would be low. Therefore, banks want to stay at the same interest rate.